DIRC Operations Weekly Report

Period: Saturday 2006/07/22 - Friday 2006/07/28

Oncall: Rolf Andreassen, Bradley Wogsland
Fast Monitoring: Bradley Wogsland, Rolf Andreassen

File location: $BFROOT/www/Detector/DIRC/WeeklyReports/report_20060728.html
Plot directory: $BFROOT/www/Detector/DIRC/WeeklyReports/Plots/report_20060728

Contents


BaBar-PEP news

PEP peak lumi: 10.6*1033 (previous week: 11.2*1033)

BaBar integrated lumi: 3.11 fb-1 (previous week: 3.13 fb-1)

Run 5b PEP delivery: 80.16 fb-1

Run 5 total delivery: 139.3 fb-1 Run 5 total BaBar recorded: 134.14 fb-1 including 9.82 fb-1 off-peak data

PEP activities this week:

  • Several down periods due to heat.
  • Thursday 27th : A 'boot' was installed on a vacuum leak in PR02. This was successful in containing the leak, dropping pressure by almost two orders of magnitude.

    PEP activities planned next week :

  • Off-peak running.

    Foreseen work opportunities (PEP down periods):

    BaBar activities/issues this week:

    DIRC activities

    See DIRC logbook for details

  • Sunday 23rd : Wiener crate yellow temperature alarms - it's hot outside! Masked the alarms.
  • Sunday 23rd (and ongoing) : An intermittently noisy channel in sector 6. We will keep an eye on it and mask it if it begins to generate deadtime.
  • Friday 28th : Paged at 0450 due to non-runnable DIRC. Problem turned out to be from a power glitch. The DFB registers were locked, so that when the pilot reconfigured, the software could not read them. On the second reconfiguration it worked properly. Plots showing the fan speed are found here : Fan 1, Fan 4. It seems there was a power glitch.
  • Friday 28th : Yellow and red scaler alarms in sector 3. MCC was notified. Not dangerous yet, keeping an eye on it.

    Backgrounds

    Background summary plot

    Scheduled activities

    Pending Work

    Manpower situation

    Links